In 2013, the prices of homes sold in Amman was skewed right with mean $255,000 and standard deviation $105,000. For 2013, the minimum percentage of homes sold in Amman with selling price between $45,000 and $465,000 is A)84% B) 75% C) 99.70% D) 95% E) 88.89%


1
Expert's answer
2021-09-02T15:35:59-0400

"\mu=255000 \\

\sigma=105000"

Let X be the price of home.

"P(45000<X<465000) = P(X<465000) -P(X<45000) \\

=P(Z< \frac{465000-255000}{105000}) -P(Z< \frac{45000 -255000}{105000}) \\

=P(Z< 2) -P(Z< -2) \\

= 0.9772-0.0228 \\

= 0.9544 \\

= 95.44 \; \%"

Answer: D) 95 %